OpenStudy (chris215):

What is the average rate of change of y with respect to x over the interval [1, 5] for the function y = 4x + 2?

hartnn (hartnn):

know the formula for rate of change?

hartnn (hartnn):

\(\Large \dfrac{f(b)-f(a)}{b-a}\)

hartnn (hartnn):

a,b = 1,5 f(x) =4x+2

OpenStudy (pawanyadav):

Average change=[f(5)-f(1)] /(5-1)

OpenStudy (chris215):

oh ok yeah I didnt knw the formula thanks! I got 1/4

hartnn (hartnn):

1/4 or 4?

OpenStudy (chris215):

1/4

hartnn (hartnn):

then show your steps?

hartnn (hartnn):

f(5) = 4*5+2 = 22 f(1) = 4*1 +2 = 6 numerator = 22-6 =16 denominator = 5-1 =4

OpenStudy (chris215):

Thank you! so it's 4

hartnn (hartnn):

yes, welcome ^_^